### v1

@@ 7.1 Ethical Considerations
This study utilizes publicly available data sources to evaluate the dialectal understanding of LLMs within the Saudi context. Care was taken to ensure that no private, identifiable, or sensitive information was included in the dataset. Given that language models may produce hallucinated or inaccurate outputs—particularly in low-resource or culturally nuanced scenarios—we stress that these outputs should not be used for critical decision-making without human oversight. The evaluation framework was carefully designed to minimize cultural bias and to avoid reinforcing stereotypes. Throughout the project, we maintained a commitment to transparency and responsible AI usage, acknowledging the inherent limitations of current LLMs when handling dialectal and culturally embedded language. Ultimately, our goal is to promote the development of fair, inclusive, and culturally aware language technologies.
7.2 Limitations
While the Absher benchmark represents a significant step toward evaluating LLMs on Saudi dialects and culture, the study is subject to several limitations. First, the evaluations were conducted under constrained computational resources, which limited our ability to explore multiple model configurations or conduct repeated trials. Second, all models were assessed using a fixed prompt structure (e.g., “Answer as if you are a Saudi…”), which may have influenced model behavior. Future work could explore the impact of prompt variation and instruction tuning. Third, although human evaluation was performed on a stratified sample to ensure dataset quality, a complete manual review of all 18,564 questions was not feasible due to the scale and diversity of the benchmark. Finally, while Absher covers various regional Saudi dialects, the representation of these dialects within the dataset is not perfectly balanced; for instance, Central and Western dialects are more prevalent. That highlights the need for more diverse and evenly distributed resources to comprehensively capture all Saudi dialectal variations in future research. Future research should leverage more extensive computational infrastructure, experiment with prompt engineering strategies, and incorporate broader expert reviews to further enhance the accuracy, reliability, and cultural validity of the benchmark evaluations.
8 Conclusion and Future Work
@@
This paper introduced Absher, a novel benchmark designed to evaluate the capabilities of large language models (LLMs) in understanding the diverse dialects and rich cultural expressions across Saudi Arabia. Unlike existing Arabic benchmarks that center on Modern Standard Arabic (MSA) or treat dialects at a coarse level, Absher offers fine-grained, region-specific evaluation through over 18,000 carefully constructed multiple-choice questions. These questions span six task types—ranging from meaning identification to cultural interpretation—and are grounded in authentic dialectal content from various Saudi regions.
@@
Looking forward, future work will expand Absher to include additional dialects, question formats, and model types. We also plan to explore few-shot and instruction-tuned settings to assess model adaptability in low-resource scenarios. Integrating sociolinguistic features and region-specific datasets may further improve alignment with real-world language use. Through these efforts, we aim to promote more inclusive, robust, and context-aware Arabic NLP systems.
@@ Appendix B Annotation Guideline
Figure 8: The interface designed using Google Sheets, which was shared with annotators to evaluate the questions.
Appendix C Detailed Performance

### v2

@@ 6.2 Ethical Considerations
This study utilizes publicly available data sources to evaluate the dialectal understanding of LLMs within the Saudi context. Care was taken to ensure that no private, identifiable, or sensitive information was included in the dataset. To guarantee this, all dialectal content—including words, phrases, and proverbs— was sourced exclusively from the publicly accessible Moajam website. This resource functions as a digital repository of aggregated dialectal vocabulary and cultural expressions. We explicitly excluded any conversational data, social media content, or personally generated user text to safeguard against the unintentional inclusion of private information or the leakage of toxic language that is often prevalent in online datasets.
The primary ethical risk in dialectal evaluation is the potential for linguistic discrimination, where underrepresented dialects are implicitly or explicitly associated with lower social status or performance failure by the model. To mitigate this critical risk, we implemented two concrete measures. First, the Absher dataset was rigorously designed to ensure fair coverage across the five major Saudi regions (Central, Western, Southern, Eastern, and Northern), deliberately challenging the common bias in NLP datasets that over-represent urban or dominant dialects. Second, the evaluation tasks focus strictly on testing linguistic competence (meaning, contextual usage, and cultural interpretation) rather than social or demographic inference, preventing the model from being scored on tasks that could encourage or propagate negative stereotypes associated with any specific regional dialect.
@@
Given that language models may produce hallucinated or inaccurate outputs—particularly in low-resource or culturally nuanced scenarios—we stress that these outputs should not be used for critical decision-making without human oversight. The evaluation framework was carefully designed to minimize cultural bias and to avoid reinforcing stereotypes.
Finally, our commitment to responsible AI usage extended to our human validation processes. The validation, which was performed by native speakers from each of the five regions to confirm dialectal authenticity, was governed. Throughout the project, we maintained a commitment to transparency and responsible AI usage, acknowledging the inherent limitations of current LLMs, with the ultimate goal of promoting the development of fair, inclusive, and culturally aware language technologies.
6.3 Conclusion and Future Work
@@
This paper introduced Absher, a novel benchmark designed to evaluate the capabilities of LLMs in understanding the diverse dialects and rich cultural expressions across Saudi Arabia. Unlike existing Arabic benchmarks that focus on Modern Standard Arabic (MSA) or treat dialects at a coarse level, Absher offers fine-grained, region-specific evaluation through over 18,000 carefully constructed multiple-choice questions. These questions span six task types—ranging from meaning identification to cultural interpretation—and are grounded in authentic dialectal content from various Saudi regions. These questions are derived from a curated dataset of dialectal words, phrases, and proverbs sourced from five major Saudi regions, and systematically span six distinct task categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition. This comprehensive design enables multifaceted evaluation across both linguistic and cultural dimensions.
@@
Looking forward, future work will expand Absher to include additional dialects, question formats, and model types. We plan to conduct more in-depth evaluations using alternative strategies, such as few-shot and fine-tuning scenarios, to better explore models’ capabilities across dialectal variations and cultural contexts. In addition, we aim to perform deeper causal analyses to understand the factors driving model performance across dialects and task types, providing clearer insights into how linguistic and cultural features influence LLM behavior. Integrating sociolinguistic features and region-specific datasets may further improve alignment with real-world language use. Through these efforts, we aim to promote more inclusive, robust, and context-aware Arabic NLP systems.
@@
CRediT authorship contribution statement
Renad Al-Monef: Methodology, Data Curation, Software, Validation, Formal analysis, Visualization, Writing - original draft. Hassan Alhuzali: Conceptualization, Supervision, Validation, Writing - review & editing. Nora Alturayeif: Supervision, Writing - review & editing. Ashwag Alasmari: Conceptualization, Supervision, Validation, Writing - review & editing.
@@ Acknowledgment
The authors would like also to extend their appreciation to the Deanship of Research and Graduate Studies at King Khalid University for funding this work through small group research under grant number RGP1/69/46. The authors also express their sincere gratitude to Al Bandari Ali Alqahtani , Zainab Abdulaziz Dalim for their invaluable support and insightful contributions during the evaluation of Absher benchmark
@@ Appendix B Annotation Guideline
Appendix C The Annotation interface
Appendix D Detailed Performance
Table 10: Detailed evaluation of the results across content and question types, with the best results highlighted in dark green to represent the highest scores.